OpenStudy (anonymous):

Write an equation of a line parallel to y = 1 over 4x − 6 in slope-intercept form that passes through the point (−1, 5). y= 1 over 4x − 1 y= 1 over 4x + 5.25 y= −4x + 5 y= −4x − 1.25

OpenStudy (anonymous):

A line parallel to y = 1/4x -6 will have the same slope as this line. So, now we just need to find the y intercept of a line with a slope of 1/4 that passes through the point (-1,5) Do you know how to find the y intercept with this information?

OpenStudy (anonymous):

no i forgot how to do it

OpenStudy (anonymous):

but i think the formula is y=mx+b i think

OpenStudy (anonymous):

Yup, so we have the formula y = mx + b m = slope b = y intercept We have the point (-1,5) and the slope 1/4 The point is saying that x = -1 and y = 5, right? So, now let's put these numbers into the formula and solve for b (y intercept) y = mx + b (5) = (1/4)(-1) + b 5 = -1/4 + b ---Add 1/4 to both sides 5 and 1/4 = b So now we know b (the y intercept) and m (the slope) so we can put this in the formula and get the final answer. y = mx + b y = 1/4x + 5 and 1/4 1/4 is the same thing as 0.25, so now we have y = 1/4x +5.25, which was your second option. :) Hope this helps.

OpenStudy (anonymous):

so the answer is B right

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

oh okay Thank You

OpenStudy (anonymous):

No problem. If it helps, here's this line graphed. You can see that it passes through (-1, 5) so we're good.

OpenStudy (anonymous):

not its okay now